### v3

# Stability of ranks under field extensions

Abstract.
This paper studies the stability of tensor ranks under field extensions. Our main contributions are fourfold: (1) We prove that the analytic rank is stable under field extensions. (2) We establish the equivalence between the partition rank vs. analytic rank conjecture and the stability conjecture for partition rank. We also prove that they are equivalent to other two important conjectures. (3) We resolve the Adiprasito-Kazhdan-Ziegler conjecture on the stability of the slice rank of linear subspaces under field extensions. (4) As an application of (1), we show that the geometric rank is equal to the analytic rank up to a constant factor.
1. Introduction
Given a tensor $T\in\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$, we denote by $\operatorname{PR}_{\mathbb{K}}(T)$ (resp. $\operatorname{SR}_{\mathbb{K}}(T)$, $\operatorname{AR}_{\mathbb{K}}(T)$, $\operatorname{GR}_{\mathbb{K}}(T)$) the partition rank [^1] It is also called the Schmidt rank in number theory [37]. [36] (resp. slice rank [43], analytic rank [21], geometric rank [28]) of $T$ (cf.  (3)–(6)). These ranks were first introduced to study various problems in additive combinatorics [22, 23, 21, 43, 13, 16, 36], but very soon they also received great attention from other fields such as algebraic geometry [27, 18], coding theory [26, 4] and algebraic complexity theory [5, 28]. Arguably, the most fundamental problems concerning these ranks are their relations to each other and their stability under field extensions. The stability of a function $f$ of tensors under field extensions means the existence of two constants $c\coloneqq c(d,\mathbb{K})>0$ and $C\coloneqq C(d,\mathbb{K})>0$ such that for any $T\in\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$ and any (finite) algebraic extension $\mathbb{F}/\mathbb{K}$, we have
$cf(T)\leq f\left(T^{\mathbb{F}}\right)\leq Cf(T),$
where $T^{\mathbb{F}}$ is the tensor in $\mathbb{F}^{n_{1}}\otimes\cdots\otimes\mathbb{F}^{n_{d}}$ determined by $T$. Here and in the rest of the paper, we use $C(d,\mathbb{K})$ to indicate the dependence of the constant $C$ on $d$ and $\mathbb{K}$. We remark that in the context of algebraic geometry [24, 12, 31], the stability of $f$ may be rephrased as the invariance of the order of $f$ under base changes.
Although the stability of ranks plays an essential role in existing works [6, 10, 11, 1, 34, 14, 9], there is no systematic investigation of the stability in the literature, to the best of our knowledge. The goal of this paper is to study the stability of aforementioned tensor ranks, and establish relations among them by the stability. To better illustrate our contributions, we first recall the following three well-known conjectures.
Conjecture 1.1 (Partition rank vs. analytic rank conjecture).
[21, 33, 1] There exists a constant $C\coloneqq C(d,q)>0$ such that
$\operatorname{PR}_{\mathbb{F}_{q}}(T)\leq C\operatorname{AR}_{\mathbb{F}_{q}}(T)$
for any $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$.
Conjecture 1.2 (Partition rank vs. geometric rank conjecture).
[37, 1, 9] There exists a constant $C\coloneqq C(d,q)>0$ such that
$\operatorname{PR}_{\mathbb{F}_{q}}(T)\leq C\operatorname{GR}_{\mathbb{F}_{q}}(T)$
for any $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$.
Conjecture 1.3 (Stability conjecture for partition rank).
[1] There exists a constant $C\coloneqq C(d,q)>0$ such that
$\operatorname{PR}_{\mathbb{F}_{q}}(T)\leq C\operatorname{PR}_{\overline{\mathbb{F}}_{q}}\left(T^{\overline{\mathbb{F}}_{q}}\right)$
for any $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$.
We remark that the original formulations of Conjectures 1.11.3 actually require the constant $C$ to be only dependent on $d$. However, for instance, it was proved in [9] that the original version of Conjecture 1.1 is true over finite fields of sufficiently large cardinality depending on the analytic rank. Thus, to completely resolve the original version of Conjecture 1.1, one needs to remove the dependence of the result in [9] on the analytic rank and prove Conjecture 1.1 for fields of small cardinalities. We also want to mention that according to [42, Remark 4], algebraic problems over fields of small cardinalities are usually considered more difficult than those over sufficiently large fields.
Main results
According to [1], one can prove that Conjecture 1.3 implies Conjecture 1.1 with an extra assumption. Only very recently, it was shown that this assumption can actually be removed [9]. In Theorem 5.6, we will prove that Conjectures 1.11.3 are all equivalent. In fact, Theorem 5.6 proves that Conjectures 1.11.3 are all equivalent to
Conjecture 1.4 (Asymptotic direct sum conjecture for partition rank).
There exists a constant $C\coloneqq C(d,q)>0$ such that
$\operatorname{PR}_{\mathbb{F}_{q}}(T)\leq C\limsup_{n\to\infty}\frac{\operatorname{PR}_{\mathbb{F}_{q}}\left(T^{\oplus n}\right)}{n}$
for any $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$.
It is noticeable that Conjecture 1.4 is a weaker version of the direct sum conjecture for partition rank [20], which is an analogue of the direct sum conjecture for cp-rank [41, 25, 29], disproved by [38]. In particular, Conjecture 1.4 is true for slice rank [20]. Thus, Theorem 5.6 trivially implies that Conjectures 1.11.3 are true for $d=3$, which were already known by results in [8, 1, 14].
One of the main ingredients in the proof of Theorem 5.6 is the stability of the analytic rank, proved in Theorem 4.2. In earlier works such as [34, Proposition 8.2] and [9, Proof of Corollary 1], it was proved that for any tensor $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$ and integers $l\geq 1,d\geq 2$,
$\operatorname{AR}_{\mathbb{F}_{q^{l}}}\left(T^{\mathbb{F}_{q^{l}}}\right)\leq l^{d-1}\operatorname{AR}_{\mathbb{F}_{q}}(T).$
Theorem 4.2 not only improves the coefficient $l^{d-1}$ to a constant depending only on $d$ and $q$, but it also shows that $\operatorname{AR}_{\mathbb{F}_{q^{l}}}\left(T^{\mathbb{F}_{q^{l}}}\right)$ is bounded below by $\operatorname{AR}_{\mathbb{F}_{q}}(T)$, up to a constant independent to $l$. As an interesting application of Theorem 4.2, we prove in Proposition 4.4 that $\operatorname{AR}_{\mathbb{F}_{q}}(T)=\operatorname{GR}_{\mathbb{F}_{q}}(T)$ up to a constant factor. Previously, it was only known that $\operatorname{GR}_{\mathbb{F}_{q}}(T)$ is bounded above by a constant multiple of $\operatorname{AR}_{\mathbb{F}_{q}}(T)$, see [8, Theorem 1] and [1, Proof of Theorem 1.13 (1)]. Moreover, we show that $\lim_{l\to\infty}\operatorname{AR}_{\mathbb{F}_{q^{l}}}\left(T^{\mathbb{F}_{q^{l}}}\right)=\operatorname{GR}_{\mathbb{F}_{q}}(T)$ in Theorem 4.3, which is an analogue of [28, Theorem 8.1].
As a corollary of the proof of Theorem 5.6, we obtain Proposition 5.3 concerning the stability of slice rank. Although a stronger stability result [14] may be obtained by geometric invariant theory, our proof of Proposition 5.3 is much more elementary. The slice rank $\operatorname{SR}_{\mathbb{K}}(\mathsf{W})$ (cf.  (22)) of a linear subspace $\mathsf{W}\subseteq\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$ was introduced in [1, Definition 1.3] and it was conjectured that $\operatorname{SR}_{\mathbb{K}}(\mathsf{W})$ is also stable under field extensions.
Conjecture 1.5.
[1, Conjecture 1.16] There is a constant $C\coloneqq C(d)>0$ such that for any field $\mathbb{K}$ and linear subspace $\mathsf{W}\subseteq\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$, we have
$\operatorname{SR}_{\mathbb{K}}(\mathsf{W})\leq C\operatorname{SR}_{\overline{\mathbb{K}}}(W\otimes\overline{\mathbb{K}}).$
Conjecture 1.5 was only confirmed for $d=2$ in [1, Theorem 1.17]. In Theorem 5.10, we resolve Conjecture 1.5 with $C=d/2+1$ for all $d\geq 2$, by associating a tensor $T_{\mathsf{W}}$ of order $d+2$ to $\mathsf{W}$ such that $\operatorname{SR}_{\mathbb{K}}(T_{\mathsf{W}})=\operatorname{SR}_{\mathbb{K}}(\mathsf{W})$.
To conclude this section, we summarize the main results of this paper in Figure 1, where $T$ is a tensor and $\mathsf{W}$ is a linear subspace of tensors, $\mathbb{F}$ is a finite extension of $\mathbb{F}_{q}$. The symbol $\approx$” means equal up to a constant factor, “?absent?$\approx$?≈ ?” means $\approx$” is conjectured to hold and all conjectures around $\operatorname{PR}_{\mathbb{F}_{q}}(T)$ are equivalent.
[…]
Figure 1. Summary of results
2. Notations
We first recall three basic operations. Given a permutation $\pi\in\mathfrak{S}_{d}$, an integer $1\leq p\leq n$ and tensors $T=u_{1}\otimes\cdots\otimes u_{d}\in\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$, $S=v_{1}\otimes\cdots\otimes v_{d}\in\mathbb{K}^{m_{1}}\otimes\cdots\otimes\mathbb{K}^{m_{d}}$, $F=f_{1}\otimes\cdots\otimes f_{p}\in(\mathbb{K}^{n_{\pi(1)}})^{\ast}\otimes\cdots\otimes(\mathbb{K}^{n_{\pi(p)}})^{\ast}$, the Kronecker product of $T$ and $S$ is
$T\boxtimes S=(u_{1}\otimes v_{1})\otimes\cdots\otimes(u_{d}\otimes v_{d})\in\mathbb{K}^{n_{1}m_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}m_{d}}.$
We denote by $\langle T,F\rangle$ or $\langle F,T\rangle$ the tensor in $\mathbb{K}^{n_{\pi(p+1)}}\otimes\cdots\otimes\mathbb{K}^{\pi(n)}$ obtained by contracting $T$ with $S$:
$\langle T,S\rangle\coloneqq\left(\prod_{i=1}^{p}f_{i}(u_{\pi(i)})\right)u_{\pi(p+1)}\otimes\cdots\otimes u_{\pi(n)}.$
The direct sum of $T$ and $S$ is
$T\oplus S\coloneqq(u_{1}\oplus v_{1})\otimes\cdots\otimes(u_{d}\oplus v_{d})\in\mathbb{F}_{q}^{n_{1}+m_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}+m_{d}}.$
We extend these operations for general tensors by linearity.
2.1. Ranks of tensors
For $S\in\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$, if there exist some partition $\pi\in\mathfrak{S}_{d}$, integer $1\leq p\leq d$, tensors $S_{1}\in\mathbb{K}^{n_{\pi(1)}}\otimes\cdots\otimes\mathbb{K}^{n_{\pi(p)}}$ and $S_{2}\in\mathbb{K}^{n_{\pi(p+1)}}\otimes\cdots\otimes\mathbb{K}^{n_{\pi(d)}}$ such that
$S=S_{1}\otimes S_{2},$
then we say $S$ has partition rank one. In particular, if $p=1$, we say $S$ has slice rank one. Here the equality is understood via the natural isomorphism
$\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}\simeq\mathbb{K}^{n_{\pi(1)}}\otimes\cdots\otimes\mathbb{K}^{n_{\pi(p)}}\otimes\mathbb{K}^{n_{\pi(p+1)}}\otimes\cdots\otimes\mathbb{K}^{n_{\pi(d)}}.$
Given $T\in\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$, we denote
$\operatorname{PR}_{\mathbb{K}}(T)$ $\coloneqq\min\left\{r\in\mathbb{N}:T=\sum_{j=1}^{r}T_{j},\;T_{j}\text{~{}has partition rank one}\right\},$ (1)
$\operatorname{SR}_{\mathbb{K}}(T)$ $\coloneqq\min\left\{r\in\mathbb{N}:T=\sum_{j=1}^{r}T_{j},\;T_{j}\text{~{}has slice rank one}\right\}.$ (2)
If $S\in\mathbb{K}^{m_{1}}\otimes\cdots\otimes\mathbb{K}^{m_{d}}$ can be written as $S=(g_{1}\otimes\cdots\otimes g_{d})T$ for some $g_{i}\in\operatorname{Hom}(\mathbb{K}^{n_{1}},\mathbb{K}^{m_{1}})$, $1\leq i\leq d$, then we write $S\preceq T$. We define cp-rank and subrank of $T$ respectively by
$\operatorname{R}_{\mathbb{K}}(T)\coloneqq\min\left\{r\in\mathbb{N}:T\preceq\operatorname{Id}_{r}\right\},\quad\operatorname{Q}_{\mathbb{K}}(T)\coloneqq\max\left\{s\in\mathbb{N}:\operatorname{Id}_{s}\preceq T\right\}.$ (3)
Here $\operatorname{Id}_{n}\in\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$ denotes the identity tensor for each positive integer $n\leq\min\{n_{j}\}_{j=1}^{d}$.
The geometric rank of $T$ is
$\operatorname{GR}_{\mathbb{K}}(T)\coloneqq\sum_{i=1}^{d}n_{i}-n_{k}-\dim Z_{k}\left(T\right),$ (4)
where
$Z_{k}\left(T\right)\coloneqq\left\{(f_{1},\dots,f_{d-1})\in\bigoplus_{1\leq i\neq k\leq d}\left(\mathbb{F}_{q}^{n_{i}}\right)^{\ast}:\langle T,f_{1}\otimes\cdots\otimes f_{d-1}\rangle=0\right\}.$ (5)
If $\mathbb{K}$ is a finite field, we also define the analytic rank of $T$:
$\operatorname{AR}_{\mathbb{K}}(T)\coloneqq-\log_{|\mathbb{K}|}\sum_{(f_{1},\dots,f_{d})\in\bigoplus_{1\leq i\leq d}\left(\mathbb{F}_{q}^{n_{i}}\right)^{\ast}}\chi(\langle T,f_{1}\otimes\cdots\otimes f_{d}\rangle),$ (6)
where $\chi:\mathbb{K}\to\mathbb{C}$ is any non-trivial additive character of $\mathbb{K}$.
2.2. Properties of ranks
For ease of reference, we collect some basic properties of ranks defined in (1)–(6).
Lemma 2.1 (Additivity of slice rank).
[20] For any $T\in\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$ and $S\in\mathbb{K}^{m_{1}}\otimes\cdots\otimes\mathbb{K}^{m_{d}}$, we have $\operatorname{SR}_{\mathbb{K}}(T\oplus S)=\operatorname{SR}_{\mathbb{K}}(T)+\operatorname{SR}_{\mathbb{K}}(S)$.
Lemma 2.2.
[32] For any $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$ and $1\leq k\leq d$, we have
$\operatorname{AR}_{\mathbb{F}_{q}}(T)=\sum_{i=1}^{d}n_{i}-n_{k}-\log_{q}\left(\left\lvert Z_{k}\left(T\right)\right\rvert\right).$
Here $Z_{k}\left(T\right)$ is defined by (5).
Lemma 2.3 (Additivity of analytic rank).
[32] For any $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$ and $S\in\mathbb{F}_{q}^{m_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{m_{d}}$, we have $\operatorname{AR}_{\mathbb{F}_{q}}(T\oplus S)=\operatorname{AR}_{\mathbb{F}_{q}}(T)+\operatorname{AR}_{\mathbb{F}_{q}}(S)$.
3. Rank and subrank of a field extension
Let $\mathbb{F}$ be a finite extension of $\mathbb{K}$. For an integer $d\geq 2$, we consider the map
$\varphi_{d}:\underbrace{\mathbb{F}\times\cdots\times\mathbb{F}}_{(d-1)\text{~{}copies}}\to\mathbb{F},\quad\varphi_{d}(x_{1},\dots,x_{d-1})=x_{1}\cdots x_{d-1}$ (7)
and denote the tensor corresponding to $\varphi_{d}$ by $T_{d,\mathbb{F}}\in\mathbb{F}^{\otimes d}$ where $\mathbb{F}^{\otimes d}\coloneqq\underbrace{\mathbb{F}\otimes\cdots\otimes\mathbb{F}}_{d\text{~{}copies}}$. The main purpose of this section is to estimate $R_{\mathbb{K}}(T_{d,\mathbb{F}})$ and $Q_{\mathbb{K}}(T_{d,\mathbb{F}})$ defined by (3).
Proposition 3.1 (Rank and subrank of field extension I).
Suppose $\mathbb{K}$ is an infinite perfect field. For any integer $d\geq 2$ and any finite extension $\mathbb{F}/\mathbb{K}$, we have
$R_{\mathbb{K}}(T_{d,\mathbb{F}})\leq(d-1)(l-1)+1,\quad Q_{\mathbb{K}}(T_{d,\mathbb{F}})\geq\left\lceil\dfrac{l-1}{d-1}\right\rceil,$
where $l\coloneqq[\mathbb{F}:\mathbb{K}]$.
Proof.
Since $[\mathbb{F}:\mathbb{K}]=l<\infty$ and $\mathbb{K}$ is perfect, we have $\mathbb{F}=\mathbb{K}[\alpha]$ for some $\alpha\in\mathbb{F}$. Thus, we obtain the following commutative diagram:
${{\underbrace{\mathbb{K}[x]_{\leq l-1}\times\cdots\times\mathbb{K}[x]_{\leq l-1}}_{(d-1)\text{~{}copies}}}}$${{\mathbb{K}[x]_{\leq(d-1)(l-1)}}}$${{\underbrace{\mathbb{K}[\alpha]\times\cdots\times\mathbb{K}[\alpha]}_{(d-1)\text{~{}copies}}}}$${{\mathbb{K}[\alpha]}}$$\scriptstyle{\Phi_{d}}$$\scriptstyle{q^{-1}\times\cdots\times q^{-1}}$$\scriptstyle{q}$$\scriptstyle{\varphi_{d}}$ (8)
where $\Phi_{d}$ is the map multiplying $(d-1)$ polynomials of degree at most $(l-1)$ and $q$ is the restriction of the quotient map $\mathbb{K}[x]\to\mathbb{K}[\alpha]$. This implies that $R_{\mathbb{K}}(T_{d,\mathbb{F}})\leq R_{\mathbb{K}}(M_{d})$ where $M_{d}\in(\mathbb{K}^{l})^{\otimes(d-1)}\otimes\mathbb{K}^{(d-1)(l-1)+1}$ is the tensor corresponding to $\Phi_{d}$. Hence it is sufficient to estimate $R_{\mathbb{K}}(M_{d})$. To that end, we denote $N\coloneqq(d-1)(l-1)$ and pick distinct elements $a_{1},\dots,a_{N+1}$ in $\mathbb{K}$. For each $11\leq j\leq N+$ 1 and $1\leq k\leq d$, we define
$\eta_{jk}\coloneqq\begin{cases}\sum_{i=0}^{l-1}a_{j}^{i}\varepsilon_{i}&\text{~{}if~{}}1\leq k\leq d-1\\ \sum_{i=0}^{N}(V^{-1})_{ji}\varepsilon_{i}&\text{~{}if~{}}k=d\\ \end{cases}$
where $(V^{-1})_{ji}$ is the $(j,i)$-th element of the inverse of the Vandermonde matrix
$V=\begin{bmatrix}1&a_{1}&\cdots&a_{1}^{N}\\ \vdots&\vdots&\ddots&\vdots\\ 1&a_{N+1}&\cdots&a_{N+1}^{N}\\ \end{bmatrix}\in\mathbb{K}^{(N+1)\times(N+1)}$
and $\varepsilon_{s}:\mathbb{K}[x]\to\mathbb{K}$ is the linear function such that $\varepsilon_{s}(x^{t})=\delta_{st}$, $0\leq s,t$. We consider
$M^{\prime}_{d}\coloneqq\sum_{j=1}^{N+1}\eta_{j1}\otimes\cdots\otimes\eta_{jd}.$
Given $f_{1},\dots,f_{d-1}\in\mathbb{K}[x]_{\leq l-1}$, it is clear that $M^{\prime}_{d}(f_{1},\dots,f_{d-1})$ computes the product polynomial $\prod_{k=1}^{d}f_{k}\in\mathbb{K}[x]_{\leq N}$ by interpolation at points $a_{1},\dots,a_{N+1}\in\mathbb{K}$. Thus $R_{\mathbb{K}}(M_{d})=R_{\mathbb{K}}(M^{\prime}_{d})\leq N+1$.
For $Q_{\mathbb{K}}(T_{d,\mathbb{F}})$, we let $m\coloneqq\lfloor(l-1)/(d-1)\rfloor$ and consider the commutative diagram:
${{\underbrace{\mathbb{K}[\alpha]\times\cdots\times\mathbb{K}[\alpha]}_{(d-1)\text{~{}copies}}}}$${{\mathbb{K}[\alpha]}}$${{\underbrace{\mathbb{K}[x]_{\leq m}\times\cdots\times\mathbb{K}[x]_{\leq m}}_{(d-1)\text{~{}copies}}}}$${{\mathbb{K}[x]_{\leq l-1}}}$$\scriptstyle{\varphi_{d}}$$\scriptstyle{q^{-1}}$$\scriptstyle{q\times\cdots\times q}$$\scriptstyle{\Psi_{d}}$ (9)
where $\Psi_{d}$ denotes the map multiplying $(d-1)$ polynomials of degree at most $m$. This implies that $Q_{\mathbb{K}}(T_{d,\mathbb{F}})\geq Q_{\mathbb{K}}(\Psi_{d})$. For distinct $b_{1},\cdots,b_{m+1}\in\mathbb{K}$, we define a linear map
$S:\mathbb{K}[x]\to\mathbb{K}^{m+1},\quad S(f)\coloneqq(f(b_{1}),\cdots,f(b_{m+1})).$
By interpolation, it is straightforward to verify that $S\mid_{\mathbb{K}[x]_{\leq m}}$ is bijective. This leads to
[…]
where $\psi_{d}$ is defined as $\psi_{d}(x_{1},\dots,x_{d-1})\coloneqq x_{1}\cdots x_{d-1}$ for the trivial $\mathbb{K}$-algebra $\mathbb{K}^{m+1}$. Since the tensor corresponding to $\psi_{d}$ is the identity tensor $\operatorname{Id}_{m+1}$, we obtain $Q_{\mathbb{K}}(T_{d,\mathbb{F}})\geq Q_{\mathbb{K}}(\Psi_{d})\geq m+1$. ∎
If $\mathbb{K}=\mathbb{F}_{q}$ is a finite field, we record the following estimate of $\operatorname{R}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}})$ for easy reference.
Theorem 3.2 (Rank of field extension II).
[3, Theorem 3] Given any integer $d>0$ and prime power $q$, there is a constant $C\coloneqq C(d,q)>0$ such that for any finite field extension $\mathbb{F}/\mathbb{F}_{q}$, we have $\operatorname{R}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}})\leq Cl$ where $l\coloneqq[\mathbb{F}:\mathbb{F}_{q}]$.
We remark that if $d=3$, Theorem 3.2 reduces to a classical result in algebraic complexity theory [30, Theorem 2]. Next we consider $\operatorname{Q}_{\mathbb{K}}(T_{d,\mathbb{F}})$ when $\mathbb{K}$ is a finite field. We borrow the idea behind the Chudnovsky-Chudnovsky method [7, 39, 2, 3] designed to obtain a sharp upper bound of $\operatorname{R}_{\mathbb{K}}(T_{3,\mathbb{F}})$. To achieve the goal, we need to establish some technical lemmas.
Lemma 3.3.
Let $\mathbb{F}/\mathbb{F}_{q}$ be an algebraic function field with genus $g$ and let $N,d,n$ be positive integers such that $N\geq g+1$, $(d-1)(N+g-1)<n$. If $\mathbb{F}/\mathbb{F}_{q}$ has at least $N$ degree one places $\mathfrak{m}_{1},\dots,\mathfrak{m}_{N}$ and a degree $n$ place $\mathfrak{m}$, then there exists a divisor $D$ such that
  1. (a)
    $v_{\mathfrak{m}}(D)=v_{\mathfrak{m}_{1}}(D)=\cdots=v_{\mathfrak{m}_{N}}(D)=0$.
  2. (b)
    The map
    $E:\mathscr{L}(D)\to\mathbb{F}_{q}^{N},\quad E(f)=(f(\mathfrak{m}_{1}),\cdots,f(\mathfrak{m}_{N}))$ (10)
    is bijective.
  3. (c)
    The map
    $S:\mathscr{L}((d-1)D)\to\mathcal{O}/\mathfrak{m},\quad S(f)=f(\mathfrak{m})$ (11)
    is injective.
Here $\mathcal{O}$ is the valuation ring associated to the place $\mathfrak{m}$.
Proof.
By Lemma A.5, there exists a non-special divisor $D_{0}$ with $\deg(D_{0})=g-1$. Let $D_{1}\coloneqq D_{0}+\mathfrak{m}_{1}+\cdots+\mathfrak{m}_{N}$. According to Lemma A.3, there exists $f_{0}\in\mathbb{F}$ such that $v_{\mathfrak{m}_{i}}((f_{0})+D_{1})=0$ and $v_{\mathfrak{m}}((f_{0})+D_{1})=0$ for $1\leq i\leq N$. Denote $D\coloneqq(f_{0})+D_{1}$. Then $\deg(D)=\deg(D_{1})=N+g-1\geq 2g$ by Lemma A.2. Thus, Lemma A.1 implies that $\ell(D)=\deg(D)+1-g=N$.
The map $E$ is well-defined since for each $1\leq i\leq N$, $v_{\mathfrak{m}_{i}}(D)=v_{\mathfrak{m}_{i}}((f_{0})+D_{1})=0$. Moreover, we observe that
$\operatorname{Ker}(E)=\mathscr{L}(D-\mathfrak{m}_{1}-\cdots-\mathfrak{m}_{N})=\mathscr{L}((f_{0})+D_{0}).$
Since $D_{0}$ is non-special, we have
$\dim\operatorname{Ker}(E)=\ell((f_{0})+D_{0})=\ell(D_{0})=\deg(D_{0})+1-g=0.$
Therefore, $E$ is an isomorphism.
Since $v_{\mathfrak{m}}(D)=0$, the map $S$ is also well-defined and
$\operatorname{Ker}(S)=\mathscr{L}((d-1)D-\mathfrak{m}).$
By assumption, $\deg((d-1)D-\mathfrak{m})=d(N+g-1)-n<0$. Hence $\ell((d-1)D-\mathfrak{m})=0$ by Lemma A.4. ∎
Lemma 3.4.
Let $n,d,N$ be positive integers. Suppose $\mathbb{F}/\mathbb{F}_{q}$ is an algebraic function field of genus $g$ such that
  1. (a)
    $N\geq g+1$ and $(d-1)(N+g-1)<n$.
  2. (b)
    $\mathbb{F}/\mathbb{F}_{q}$ has at least $N$ degree one places and one degree $n$ place.
Then $\operatorname{Q}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}_{q^{n}}})\geq N$.
Proof.
Let $\mathfrak{m}_{1},\dots,\mathfrak{m}_{N}$ be distinct degree one places and let $\mathfrak{m}$ be a degree $n$ place in Lemma 3.3. We denote by $\mathcal{O}$ (resp. $\mathcal{O}_{1},\dots,\mathcal{O}_{N}$) the discrete valuation ring determined by $\mathfrak{m}$ (resp. $\mathfrak{m}_{1},\dots,\mathfrak{m}_{N}$). By Lemma 3.3, there exists a divisor $D$ such that $\ell(D)=N$ and the $\mathbb{F}_{q}$-linear map $E$ (resp. $S$) defined in (10) (resp. (11)) is bijective (resp. injective). Since $S$ is injective, there is a $\mathbb{F}_{q}$-linear map $S^{-1}:\mathcal{O}/\mathfrak{m}\to\mathscr{L}((d-1)D)$ such that $S^{-1}\circ S=\operatorname{id}_{\mathscr{L}((d-1)D)}$.
We consider the following diagram:
[…]
Here $\widetilde{S}:\mathcal{O}\to\mathcal{O}/{\mathfrak{m}}\simeq\mathbb{F}_{q^{N}}$ and $\widetilde{E}:\mathcal{O}_{1}\bigcap\cdots\bigcap\mathcal{O}_{N}\to\mathbb{F}_{q}^{N}$ are evaluation maps defined by $\widetilde{S}(f)=f(\mathfrak{m})$ and $\widetilde{E}(f)=(f(\mathfrak{m}_{1}),\dots,f(\mathfrak{m}_{N}))$, respectively. We claim that vertical maps in (12) are well-defined and the diagram commutes. This implies that
$\operatorname{Q}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}_{q^{n}}})=\operatorname{Q}_{\mathbb{F}_{q}}(T_{d,\mathcal{O}/\mathfrak{m}})\geq\operatorname{Q}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}_{q}^{N}})=N.$
It is left to prove the claim. We first verify that vertical maps are well-defined. According to Lemma 3.3, we have $v_{\mathfrak{m}}(D)=v_{\mathfrak{m}_{1}}(D)=\cdots=v_{\mathfrak{m}_{N}}(D)$. This implies that $v_{\mathfrak{m}}(f)\geq 0$ if $f\in\mathcal{L}(D)$ and $v_{\mathfrak{m}_{j}}(f)\geq 0$ for all $1\leq j\leq N$ if $f\in\mathcal{L}((d-1)D)$, which lead to $\mathcal{L}(D)\subseteq\mathcal{O}$ and $\mathcal{L}((d-1)D)\subseteq\bigcap_{j=1}^{N}\mathcal{O}_{j}$. Therefore, maps $\widetilde{S}\circ E^{-1}$ and $\widetilde{E}\circ S^{-1}$ are well-defined.
Next, we prove the commutativity of (12). For any $a_{1},\dots,a_{d-1}\in\mathbb{F}_{q}^{N}$, we denote $f_{j}\coloneqq E^{-1}(a_{j})\in\mathscr{L}(D)$, $11\leq j\leq d-$ 1. Thus
$\psi_{d}(\widetilde{S}(f_{1}),\dots,\widetilde{S}(f_{d-1})))=\prod_{j=1}^{d-1}\widetilde{S}(f_{j})=\widetilde{S}\left(\prod_{j=1}^{d-1}f_{j}\right).$
Here the second equality holds since $\widetilde{S}$ is an $\mathbb{F}_{q}$-algebra homomorphism. We observe that $\prod_{j=1}^{d-1}f_{j}\in\mathcal{L}((d-1)D)$. This implies that
$(\widetilde{E}\circ S^{-1})\circ\psi_{d}\circ(\widetilde{S}\circ E^{-1})(a_{1},\dots,a_{d-1})=\widetilde{E}\left(\prod_{j=1}^{d-1}f_{j}\right).$
Since $\widetilde{E}$ is an $\mathbb{F}_{q}$-algebra homomorphism, we have,
$\widetilde{E}\left(\prod_{j=1}^{d-1}f_{j}\right)=\prod_{j=1}^{d-1}\widetilde{E}(f_{j})=\prod_{j=1}^{d-1}a_{i}=\varphi_{d}(a_{1},\dots,a_{d-1}).\qed$
Lemma 3.5.
If $(l-1)\geq(d-1)(n-1)$, then we have $\operatorname{Q}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}_{q^{n}}})\leq\operatorname{Q}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}_{q^{l}}})$.
Proof.
Since $\mathbb{F}_{q}$ is a perfect field, we observe that (8) (resp. (9)) holds for $(\mathbb{K},\mathbb{F})=(\mathbb{F}_{q},\mathbb{F}_{q^{n}})$ (resp. $(\mathbb{K},\mathbb{F})=(\mathbb{F}_{q},\mathbb{F}_{q^{l}})$). Thus, we have $T_{d,\mathbb{F}_{q^{n}}}\preceq M_{d,n,\mathbb{F}_{q}}\preceq T_{d,\mathbb{F}_{q^{l}}}$, from which we obtain $Q_{\mathbb{F}_{q}}(T_{d,\mathbb{F}_{q^{n}}})\leq Q_{\mathbb{F}_{q}}(T_{d,\mathbb{F}_{q^{l}}})$. Here $M_{d,n,\mathbb{F}_{q}}$ is the tensor corresponding to the multilinear map multiplying $(d-1)$ $\mathbb{F}_{q}$-polynomials of degree at most $(n-1)$:
$\Phi_{d}:\underbrace{\mathbb{F}_{q}[x]_{\leq n-1}\times\cdots\times\mathbb{F}_{q}[x]_{\leq n-1}}_{(d-1)\text{~{}copies}}\to\mathbb{F}_{q}[x]_{(d-1)(n-1)},\quad\Phi_{d}(f_{1},\dots,f_{d-1})=\prod_{j=1}^{d-1}f_{j}.\qed$
Lemma 3.6.
$\operatorname{Q}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}_{q^{mn}}})\geq\operatorname{Q}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}_{q^{m}}})\operatorname{Q}_{\mathbb{F}_{q^{m}}}(T_{d,\mathbb{F}_{q^{mn}}})$.
Proof.
We notice that
$T_{d,\mathbb{F}_{q^{m}}}\in\underbrace{\mathbb{F}_{q^{m}}\otimes_{\mathbb{F}_{q}}\cdots\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q^{m}}}_{d\text{~{}copies}},\quad T_{d,\mathbb{F}_{q^{mn}}}\in\underbrace{\mathbb{F}_{q^{mn}}\otimes_{\mathbb{F}_{q^{m}}}\cdots\otimes_{\mathbb{F}_{q^{m}}}\mathbb{F}_{q^{mn}}}_{d\text{~{}copies}}$
and
$T_{d,\mathbb{F}_{q^{m}}}\boxtimes_{\mathbb{F}^{m}_{q}}T_{d,\mathbb{F}_{q^{mn}}}=T_{d,\mathbb{F}_{q^{mn}}}\in\underbrace{\mathbb{F}_{q^{mn}}\otimes_{\mathbb{F}_{q}}\cdots\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q^{mn}}}_{d\text{~{}copies}},$
The desired inequality follows immediately from the definition of the subrank. ∎
Theorem 3.7 (Subrank of field extension II).
For any positive integers $d,n$ and prime power $q$, we have $\operatorname{Q}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}_{q^{n}}})\geq\frac{n}{48d^{2}q}$.
Proof.
We denote $\sigma_{q}(n)\coloneqq Q_{\mathbb{F}_{q}}(T_{d,\mathbb{F}_{q^{n}}})$ and we first estimate $\sigma_{q^{2}}(n)$. For each positive integer $k$, we let $g_{k}$ and $N_{k}$ (if $k\geq 3$) be positive integers defined as in Lemma A.6. We consider
$m\coloneqq\max\{k\in\mathbb{N}:n\geq 2dg_{k}+3\}$
and for $m\geq 3$
$N\coloneqq\max\{N^{\prime}\in\mathbb{N}:n>d(N^{\prime}+g_{m}-1),\;N_{m}\geq N^{\prime}\}.$
If $n\geq 2dg_{3}+3$, then $m\geq 3$. By Lemma A.6, there exists an algebraic field extension $\mathbb{F}/\mathbb{F}_{q^{2}}$ of genus $g_{m}$ admitting at least $N_{m}$ degree one places. Since $n\geq 2dg_{m}+3\geq 4g_{m}+3$, we have $N\geq g_{m}+1$ and Lemma A.7 implies that $\mathbb{F}/\mathbb{F}_{q^{2}}$ admits a place of degree $n$. Lemma 3.4 ensures that
$\sigma_{q^{2}}(n)\geq N.$
By (30), we have $N_{m}\geq g_{m}$. The maximality of $m$ and $N$ leads to
$2dg_{m+1}+3>n,\quad N+1>\min\left\{\frac{n}{d}-g_{m},N_{m}\right\}\geq\min\{g_{m},N_{m}\}=g_{m}$
According to (29), we obtain by a direct calculation that $g_{m+1}<2qg_{m}$, from which we have
$N>g_{m}-1>\frac{g_{m+1}}{2q}-1>\frac{n-3}{4qd}-1\geq\frac{n}{12qd}.$
For $n<2dg_{3}+3\leq 2dq^{3}$, we split the discussion into two cases. If $q^{2}\geq n/d$, Proposition 3.1 implies that $\sigma_{q^{2}}(n)\geq n/d$. If $q^{2}<n/d$, a similar interpolation argument as in the proof of Proposition 3.1 shows $\sigma_{q^{2}}(n)\geq q^{2}$. Therefore, we obtain $\sigma_{q^{2}}(n)\geq\min(n/d,q^{2})\geq\frac{1}{2qd}n$. To complete the proof, we recall from Lemma 3.6 that
$\sigma_{q}(2n)\geq\sigma_{q}(2)\sigma_{q^{2}}(n)\geq\frac{n}{12qd}.$
Given an odd integer $l>2d$, we let $n\coloneqq\left\lfloor\frac{1}{2}\left(\frac{l-1}{d-1}+1\right)\right\rfloor$ so that
$l-1\geq(d-1)(2n-1),\quad n\geq\frac{1}{2}\left(\frac{l-1}{d-1}-1\right)>\frac{l}{4d}.$
Lemma 3.5 implies $\sigma_{q}(l)\geq\sigma_{q}(2n)\geq n/12qd>l/48qd^{2}$ and this completes the proof. ∎
4. Analytic rank and Geometric rank under field extension
This section is devoted to a discussion of the behaviour of $\operatorname{AR}_{\mathbb{K}}(T)$ and $\operatorname{GR}_{\mathbb{K}}(T)$ under field extensions, defined respectively in (6) and (4). For each $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$ and a field extension $\mathbb{F}/\mathbb{F}_{q}$, we denote by $T^{\mathbb{F}}$ the tensor in $\mathbb{F}^{n_{1}}\otimes\cdots\otimes\mathbb{F}^{n_{d}}$ determined by $T$. Let $T_{d,\mathbb{F}}$ be the tensor in $\mathbb{F}^{\otimes d}$ corresponding to the multiplication map $\varphi_{d}$ defined in (7). By definition, $T\boxtimes T_{d,\mathbb{F}}$ is a tensor over $\mathbb{F}_{q}$ and we have
$T^{\mathbb{F}}=T\boxtimes T_{d,\mathbb{F}}.$
Here $T^{\mathbb{F}}$ is regarded as a tensor over $\mathbb{F}_{q}$. Moreover, we remark that $T^{\mathbb{F}}\neq T$ as tensors over $\mathbb{F}_{q}$.
Lemma 4.1.
Given $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$ and $S\in\mathbb{F}_{q}^{m_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{m_{d}}$, if there exist $M_{i}\in\mathbb{F}_{q}^{n_{i}\times m_{i}}$ for each $1\leq i\leq d$ such that $(M_{1}\otimes\cdots\otimes M_{d})S=T$, then $\operatorname{AR}_{\mathbb{F}_{q}}(T)\leq\operatorname{AR}_{\mathbb{F}_{q}}(S)$.
Proof.
Since $M_{1}\otimes\cdots\otimes M_{d}=(M_{1}\otimes I_{m_{2}}\otimes\cdots\otimes I_{m_{d}})(I_{n_{1}}\otimes M_{2}\otimes\cdots\otimes I_{m_{d}})\cdots(I_{n_{1}}\otimes I_{n_{2}}\otimes\cdots\otimes M_{d})$, it suffices to assume that $m_{i}=n_{i}$ and $M_{i}=I_{n_{i}}$ for each $11\leq i\leq d-$ 1. Denote $N\coloneqq\prod_{i=1}^{d-1}n_{i}$. Let $T^{(d)}\in\mathbb{F}_{q}^{n_{d}\times N}$ (resp. $S^{(d)}\in\mathbb{F}_{q}^{m_{d}\times N}$) be the $d$-th flattening of $T$ (resp. $S$). Then we have
$(M_{d}\otimes I_{N})S^{(d)}=T^{(d)}.$
In particular, $Z_{d}\left(S\right)\subseteq Z_{d}\left(T\right)$. By Lemma 2.2, we have
$\operatorname{AR}_{\mathbb{F}_{q}}(T)$ $=\sum_{i=1}^{d}n_{i}-n_{d}-\log_{q}\left(\left\lvert Z_{d}\left(T\right)\right\rvert\right)$
$\leq\sum_{i=1}^{d}n_{i}-n_{d}-\log_{q}\left(\left\lvert Z_{d}\left(S\right)\right\rvert\right)$
$=\operatorname{AR}_{\mathbb{F}_{q}}(S).\qed$
Theorem 4.2 (Stability of analytic/geometric rank).
Given an integer $d>0$ and prime power $q$, there exist constant numbers $c\coloneqq c(d,q)>0$ and $C\coloneqq C(d,q)>0$ such that for any positive integers $n_{1},\dots,n_{d}$ and tensor $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$, we have
$c\operatorname{AR}_{\mathbb{F}_{q}}(T)\leq\operatorname{AR}_{\mathbb{F}}\left(T^{\mathbb{F}}\right)\leq C\operatorname{AR}_{\mathbb{F}_{q}}(T)$
for any finite extension of $\mathbb{F}/\mathbb{F}_{q}$. Moreover, we also have $\operatorname{GR}_{\mathbb{F}_{q}}(T)=\operatorname{GR}_{\overline{\mathbb{F}}_{q}}\left(T^{\overline{\mathbb{F}}_{q}}\right)$.
Proof.
The equality $\operatorname{GR}_{\mathbb{F}_{q}}(T)=\operatorname{GR}_{\overline{\mathbb{F}}_{q}}\left(T^{\overline{\mathbb{F}}_{q}}\right)$ is derived from the definition of the geometric rank. Let $\mathbb{F}=\mathbb{F}_{q^{l}}$ be a finite extension of $\mathbb{F}_{q}$. By Lemma 2.2, we have
$\operatorname{AR}_{\mathbb{F}}\left(T^{\mathbb{F}}\right)$ $=\sum_{i=1}^{d}n_{i}-n_{d}-\log_{\lvert\mathbb{F}\rvert}\left(\left\lvert Z_{d}\left(T^{\mathbb{F}}\right)\right\rvert\right)$
$=\sum_{i=1}^{d}n_{i}-n_{d}-\frac{\log_{|\mathbb{F}_{q}|}(\lvert Z_{d}(T\boxtimes T_{d,\mathbb{F}})\rvert)}{l}.$
Since $\mathbb{F}_{q}^{n_{i}}\otimes\mathbb{F}\simeq\mathbb{F}_{q}^{n_{i}l}$ as $\mathbb{F}_{q}$-vector spaces, we may regard $T\boxtimes T_{d,\mathbb{F}}$ as a tensor in $\mathbb{F}_{q}^{n_{1}l}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}l}$. By Lemma 2.2 again, we have
$\log_{\lvert\mathbb{F}_{q}\rvert}\left(\lvert Z_{d}(T\boxtimes T_{d,\mathbb{F}})\rvert\right)=\sum_{i=1}^{d}n_{i}l-n_{d}l-\operatorname{AR}_{\mathbb{F}_{q}}(T\boxtimes T_{d,\mathbb{F}}).$
Thus, we obtain
$\operatorname{AR}_{\mathbb{F}}\left(T^{\mathbb{F}}\right)=\frac{\operatorname{AR}_{\mathbb{F}_{q}}(T\boxtimes T_{d,\mathbb{F}})}{l}.$ (13)
Denote $s\coloneqq\operatorname{Q}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}})$. By definition, there exist matrices $M_{i}\in\mathbb{F}_{q}^{s\times l}$ for $1\leq i\leq d$ such that $(M_{1}\otimes\cdots\otimes M_{d})T_{d,\mathbb{F}}=\operatorname{Id}_{s}$, where $\operatorname{Id}_{s}$ is the diagonal tensor in $(\mathbb{F}_{q}^{s})^{\otimes d}$. Therefore, we have
$\left((I_{n_{1}}\boxtimes M_{1})\otimes\cdots\otimes(I_{n_{d}}\boxtimes M_{d})\right)(T\boxtimes T_{d,\mathbb{F}})=T\boxtimes\operatorname{Id}_{s}=T^{\oplus s}.$
By Lemmas 4.1, 2.3 and Theorem 3.7, we have
$\operatorname{AR}_{\mathbb{F}_{q}}(T\boxtimes_{\mathbb{F}_{q}}T_{d,\mathbb{F}})\geq\operatorname{AR}_{\mathbb{F}_{q}}(T^{\oplus s})=s\operatorname{AR}_{\mathbb{F}_{q}}(T)\geq c_{d,q}l\operatorname{AR}_{\mathbb{F}_{q}}(T)$ (14)
for some constant number $c(d,q)>0$. Combining (13) and (14), we obtain
$c_{d,q}\operatorname{AR}_{\mathbb{F}_{q}}(T)\leq\operatorname{AR}_{\mathbb{F}}(T^{\mathbb{F}}).$
To conclude the proof, we replace $\operatorname{Q}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}})$ by $\operatorname{R}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}})$ in the above argument and the other inequality follows immediately from Theorem 3.2. ∎
As an application of Theorem 4.2, we will prove that the analytic rank and the geometric rank are the same up to a constant factor. To this end, we establish the following asymptotic relation between them with respect to field extensions.
Theorem 4.3 (Geometric rank = limit of analytic rank).
Given any tensor $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$, we have
$\lim_{l\to\infty}\operatorname{AR}_{\mathbb{F}_{q^{l}}}\left(T^{\mathbb{F}_{q^{l}}}\right)=\operatorname{GR}_{\mathbb{F}_{q}}(T).$
Proof.
Let $\mathbb{F}=\mathbb{F}_{q^{l}}$ be a field extension of $\mathbb{F}_{q}$. We recall from (5) that $Z_{d}\left(T\right)$ is an affine cone in $\mathbb{F}_{q}^{m}$ where $m\coloneqq\sum_{i=1}^{d-1}n_{i}$. Moreover, we have $Z_{d}\left(T^{\mathbb{F}}\right)=Z_{d}\left(T\right)(\mathbb{F})$. Let $\pi:\mathbb{F}^{m}\setminus\{0\}\to\mathbb{P}_{\mathbb{F}}^{m-1}$ be the projectivization map and let $W\coloneqq\pi\left(Z_{d}\left(T^{\mathbb{F}}\right)\setminus\{0\}\right)$. Clearly, we have
$(q^{l}-1)\lvert W\lvert+1=\left\lvert Z_{d}\left(T\right)(\mathbb{F})\right\rvert.$ (15)
According to Theorem B.1, for each irreducible component $X$ of $W$, there is a constant number $C(X)>0$ such that
$\left\lvert\lvert X(\mathbb{F})\rvert-\frac{q^{l(r+1)}-1}{q^{l}-1}\right\rvert\leq(d-1)(d-2)q^{l(r-\frac{1}{2})}+C_{X}q^{l(r-1)}=O\left(q^{l(r-\frac{1}{2})}\right),$ (16)
where $d\coloneqq\deg(X)$ and $r\coloneqq\dim(X)\leq\dim Z_{d}(T)-1=m-\operatorname{GR}_{\mathbb{F}_{q}}(T)-1$. We notice that there exits a component $X$ such that $\dim(X)=m-\operatorname{GR}_{\mathbb{F}_{q}}(T)-1$. Hence by (15) and (16) we may derive
$q^{l(m-\operatorname{GR}_{\mathbb{F}_{q}}(T))}-O\left(q^{l\left(m-\operatorname{GR}_{\mathbb{F}_{q}}(T)-\frac{1}{2}\right)}\right)\leq\lvert Z_{d}(T)(\mathbb{F})\rvert\leq Cq^{l(m-\operatorname{GR}_{\mathbb{F}_{q}}(T))}$
for some constant number $C>0$ that only depends on $T$ and $q$. Therefore, we have
$\lim_{l\to\infty}\log_{q^{l}}\lvert Z_{d}(T)(\mathbb{F})\rvert=m-\operatorname{GR}_{\mathbb{F}_{q}}(T)$
and this together with Lemma 2.2 completes the proof. ∎
It is proved in [28, Theorem 8.1] that for any $T\in\mathbb{Z}^{n_{1}}\otimes\cdots\otimes\mathbb{Z}^{n_{d}}$,
$\liminf_{p}\operatorname{AR}_{\mathbb{Z}_{p}}\left(T^{\mathbb{Z}_{p}}\right)=\operatorname{GR}_{\mathbb{Z}}(T),$ (17)
where $p$ runs through all prime numbers. We remark that both Theorem 4.3 and (17) can be recognized as: the geometric rank of a tensor is equal to the limit of its analytic ranks under base changes. Theorem 4.3 is concerned with this property for field extensions of a finite field, while (17) validates this property for the base change from $\mathbb{Z}$ to $\mathbb{Z}_{p}$.
As a direct consequence of Theorems 4.2 and 4.3, we have the proposition that follows.
Proposition 4.4 (Geometric rank $\approx$ analytic rank).
For any positive integer $d$ and prime power $q$, there are constant numbers $C\coloneqq C(d,q)>0$ and $c\coloneqq c(d,q)>0$ such that
$c\operatorname{AR}_{\mathbb{F}_{q}}(T)\leq\operatorname{GR}_{\mathbb{F}_{q}}(T)\leq C\operatorname{AR}_{\mathbb{F}_{q}}(T)$
for any $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$.
According to [1, Proof of Theorem 1.13 (1)], we have $C\leq(1-\log_{q}(d))^{-1}$. Furthermore, the proof of Theorem 4.2 indicates that $c\geq(48d^{2}q)^{-1}$. As far as we are aware, the existence of $c$ is not previously known. Let $\mathbb{F}$ be a field and let $\langle n,n,n\rangle_{\mathbb{F}}$ be the tensor in $\mathbb{F}^{n}\otimes\mathbb{F}^{n}\otimes\mathbb{F}^{n}$ corresponding to the $n\times n$ matrix multiplication. By [28, Theorem 3], we have $\operatorname{GR}_{\mathbb{F}}(\langle n,n,n\rangle)=\lceil 3n^{2}/4\rceil$. Applying Theorems 4.2, 4.3 and Proposition 4.4 to $\langle n,n,n\rangle_{\mathbb{F}_{q}}$, we obtain the following result, which may be of independent interest.
Corollary 4.5 (Analytic rank of matrix multiplication).
For any prime power $q$, we have:
  1. (a)
    For any integer $l>0$,
    $(432q)^{-1}\operatorname{AR}_{\mathbb{F}_{q}}\left(\langle n,n,n\rangle_{\mathbb{F}_{q}}\right)\leq\operatorname{AR}_{\mathbb{F}_{q^{l}}}\left(\langle n,n,n\rangle_{\mathbb{F}_{q^{l}}}\right)\leq(1-\log_{q}(3))^{-1}\operatorname{AR}_{\mathbb{F}_{q}}(\langle n,n,n\rangle_{\mathbb{F}_{q}}).$
  2. (b)
    For each $\varepsilon>0$, there exists an integer $l_{0}>0$ such that for any $l\geq l_{0}$
    $\left\lvert\operatorname{AR}_{\mathbb{F}_{q^{l}}}\left(\langle n,n,n\rangle_{\mathbb{F}_{q^{l}}}\right)-\lceil 3n^{2}/4\rceil\right\rvert\leq\varepsilon.$
  3. (c)
    $(1-\log_{q}(3))\lceil 3n^{2}/4\rceil\leq\operatorname{AR}_{\mathbb{F}_{q}}(\langle n,n,n\rangle_{\mathbb{F}_{q}})\leq 108q(3n^{2}+1)$. In particular,
    $\limsup_{n\to\infty}\log_{n}\operatorname{AR}_{\mathbb{F}_{q}}(\langle n,n,n\rangle_{\mathbb{F}_{q}})=2.$
We remark that (b) and (c) can also be obtained by a direct calculation. Indeed, it is straightforward to verify that $\operatorname{AR}_{\mathbb{F}_{q}}(\langle n,n,n\rangle_{\mathbb{F}_{q}})=\log_{q}\left\lvert\{(A,B)\in\mathbb{F}_{q}^{n\times n}\times\mathbb{F}_{q}^{n\times n}:AB=0\}\right\rvert$, from which $\operatorname{AR}_{\mathbb{F}_{q}}(\langle n,n,n\rangle_{\mathbb{F}_{q}})$ can be estimated by counting the number of $B\in\mathbb{F}_{q}^{n\times n}$ such that $AB=0$ when $\operatorname{rank}(A)=r$ for each $0\leq r\leq n$.
5. Partition rank and slice rank under field extension
In this section, we investigate the behaviour of the partition rank and the slice rank under field extensions. Let $\mathbb{K}$ be a field and let $\mathbb{F}$ be a finite extension of $\mathbb{K}$. We notice that any tensor $S\in\mathbb{F}^{n_{1}}\otimes\cdots\otimes\mathbb{F}^{n_{d}}$ can be viewed as a tensor over $\mathbb{K}$ via the $\mathbb{K}$-linear isomorphism:
$\mathbb{F}^{n_{1}}\otimes\cdots\otimes\mathbb{F}^{n_{d}}\simeq\left(\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}\right)\otimes_{\mathbb{K}}\mathbb{F}.$
Thus, both $\operatorname{PR}_{\mathbb{F}}(S)$ and $\operatorname{PR}_{\mathbb{K}}(S)$ are defined. Moreover, they are related by an inequality.
Lemma 5.1.
Suppose that $\mathbb{K}$ is either a finite field or a perfect infinite field and $\mathbb{F}$ is a finite extension of $\mathbb{K}$. Given any tensor $S\in\mathbb{F}^{n_{1}}\otimes\cdots\otimes\mathbb{F}^{n_{d}}$ with $\operatorname{PR}_{\mathbb{F}}(S)=r$, we have
$r\leq\operatorname{PR}_{\mathbb{K}}(S)\leq rR_{\mathbb{K}}(T_{3,\mathbb{F}}).$
Proof.
The first inequality follows immediately from the definition. Since $\operatorname{PR}_{\mathbb{F}}(S)=r$, we may write $S=\sum_{i=1}^{r}S_{i}$ for some $S_{i}\in\mathbb{F}^{n_{1}}\otimes\cdots\otimes\mathbb{F}^{n_{d}}$ such that $\operatorname{PR}_{\mathbb{F}}(S_{i})=1$. This implies that $\operatorname{PR}_{\mathbb{K}}(S)\leq\sum_{i=1}^{r}\operatorname{PR}_{\mathbb{K}}(S_{i})$ and it reduces to prove the lemma for $r=1$. If $\operatorname{PR}_{\mathbb{F}}(S)=1$, we may write $S=a\otimes b$ for some $a\in\mathbb{F}^{n_{\pi(1)}}\otimes\cdots\otimes\mathbb{F}^{n_{\pi(p)}}$ and $b\in\mathbb{F}^{n_{\pi(p+1)}}\otimes\cdots\otimes\mathbb{F}^{n_{\pi(d)}}$. Here $\pi\in\mathfrak{S}_{d}$ and $a\otimes b$ is identified with the tensor in $\mathbb{F}^{n_{1}}\otimes\cdots\otimes\mathbb{F}^{n_{d}}$ via the natural isomorphism
$\mathbb{F}^{n_{\pi(1)}}\otimes\cdots\otimes\mathbb{F}^{n_{\pi(d)}}\simeq\mathbb{F}^{n_{1}}\otimes\cdots\otimes\mathbb{F}^{n_{d}}.$
For convenience, we always adopt this convention in the sequel.
We observe that $\mathbb{F}^{n_{\pi(1)}}\otimes\cdots\otimes\mathbb{F}^{n_{\pi(p)}}=\left(\mathbb{K}^{n_{\pi(1)}}\otimes\cdots\otimes\mathbb{K}^{n_{\pi(p)}}\right)\otimes_{\mathbb{K}}\mathbb{F}$, from which we may identify $a$ with a $\mathbb{K}$-multilinear map $f_{a}:\mathbb{K}^{n_{\pi(1)}}\times\cdots\times\mathbb{K}^{n_{\pi(p)}}\to\mathbb{F}$. Similarly, we identify $b$ (resp. $a\otimes b$) with a $\mathbb{K}$-multilinear map $f_{b}:\mathbb{K}^{n_{\pi(p+1)}}\times\cdots\times\mathbb{K}^{n_{\pi(d)}}\to\mathbb{F}$ (resp. $f_{a\otimes b}:\mathbb{K}^{n_{1}}\times\cdots\times\mathbb{K}^{n_{d}}\to\mathbb{F}$). We notice that as $K$-multilinear maps, it holds that $f_{a\otimes b}=T_{3,\mathbb{F}}\circ(f_{a}\otimes f_{b})$. Suppose that $T_{3,\mathbb{F}}=\sum_{j=1}^{s}\alpha_{j}\otimes\beta_{j}\otimes u_{j}$ where $s=R_{\mathbb{K}}(T_{3,\mathbb{F}})$, $\alpha_{j}$ and $\beta_{j}$ are $\mathbb{K}$-linear forms for $1\leq j\leq s$. Then we can write
$f_{a\otimes b}=\sum_{j=1}^{s}(\alpha_{j}\circ f_{a})\left[(\beta_{j}\circ f_{b})u_{j}\right].$
This implies that $\operatorname{PR}_{\mathbb{K}}(a\otimes b)\leq s=R_{\mathbb{K}}(T_{3,\mathbb{F}})$. ∎
Lemma 5.2.
Suppose $\mathbb{K}$ is either a finite field or a perfect infinite field. Let $T\in\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$ be a tensor. If there exists a constant $c\coloneqq c(d,\mathbb{K})>0$ such that
$\limsup_{n\to\infty}\frac{\operatorname{PR}_{\mathbb{K}}\left(T^{\oplus n}\right)}{n}\geq c\operatorname{PR}_{\mathbb{K}}(T),$
then there exists a constant $C\coloneqq C(d,\mathbb{K})>0$ such that
$\operatorname{PR}_{\mathbb{K}}(T)\leq C\operatorname{PR}_{\overline{\mathbb{K}}}\left(T^{\overline{\mathbb{K}}}\right).$
The same is true if we replace the partition rank in the above statement by slice rank or cp-rank.
Proof.
Proofs for slice rank and cp-rank are the same as that for partition rank, thus we only prove for partition rank. Denote $r\coloneqq\operatorname{PR}_{\overline{\mathbb{K}}}\left(T^{\overline{\mathbb{K}}}\right)$. Then for each $1\leq j\leq r$, there are $\pi\coloneqq\pi_{j}\in\mathfrak{S}_{d}$, $1\leq p\coloneqq p_{j}\leq d$, $a_{j}\in\overline{\mathbb{K}}^{n_{\pi(1)}}\otimes\cdots\otimes\overline{\mathbb{K}}^{n_{\pi(p)}}$ and $b_{j}\in\overline{\mathbb{K}}^{n_{\pi(p+1)}}\otimes\cdots\otimes\overline{\mathbb{K}}^{n_{\pi(d)}}$ such that
$T^{\overline{\mathbb{K}}}=\sum_{j=1}^{r}a_{j}\otimes b_{j}.$ (18)
Since $a_{j}$’s and $b_{j}$’s are tensors over $\overline{\mathbb{K}}$, there exists a finite extension $\mathbb{F}/\mathbb{K}$ such that for each $1\leq j\leq r$,
$a_{j}\in\mathbb{F}^{n_{\pi(1)}}\otimes\cdots\otimes\mathbb{F}^{n_{\pi(p)}},\quad b_{j}\in\mathbb{F}^{n_{\pi(p+1)}}\otimes\cdots\otimes\mathbb{F}^{n_{\pi(d)}}.$ (19)
In particular, we have
$r=\operatorname{PR}_{\overline{\mathbb{K}}}\left(T^{\overline{\mathbb{K}}}\right)=\operatorname{PR}_{\mathbb{F}}\left(T^{\mathbb{F}}\right),\quad T^{\overline{\mathbb{K}}}=T^{\mathbb{F}}=T\boxtimes_{\mathbb{K}}T_{d,\mathbb{F}}.$
By Lemma 5.1, Proposition 3.1 and Theorem 3.2, there exists $c_{1}\coloneqq c_{1}(d,\mathbb{K})$ such that $\operatorname{PR}_{\mathbb{K}}(T\boxtimes_{\mathbb{K}}T_{d,\mathbb{F}})\leq c_{1}lr$. On the other side, Proposition 3.1 and Theorem 3.7 implies that there exists some constant $c_{2}\coloneqq c_{2}(d,\mathbb{K})$ such that $Q_{\mathbb{K}}(T_{d,\mathbb{F}})\geq c_{2}l$. This implies $T\boxtimes_{\mathbb{K}}T_{d,\mathbb{F}}\succeq T^{\oplus c_{2}l}$ and by the assumption we have
$c_{1}lr\geq\operatorname{PR}_{\mathbb{K}}(T\boxtimes_{\mathbb{K}}T_{d,\mathbb{F}})\geq cc_{2}l\operatorname{PR}_{\mathbb{K}}(T).$
Therefore, we obtain the desired inequality by setting $C\coloneqq c_{1}/cc_{2}$. ∎
We notice that the assumption in Lemma 5.2 is satisfied by slice rank. Indeed, according to Lemma 2.1, we may take $c=1$ for slice rank. As an application of Lemma 5.2, we derive the stability of the slice rank under field extensions.
Proposition 5.3 (Stability of slice rank).
Let $T\in\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$ be a tensor. If $\mathbb{K}$ is a perfect infinite field, then we have
$\operatorname{SR}_{\mathbb{K}}(T)\leq d^{2}\operatorname{SR}_{\overline{\mathbb{K}}}\left(T^{\overline{\mathbb{K}}}\right).$
If $\mathbb{K}=\mathbb{F}_{q}$, then there is a constant number $C\coloneqq C(q,d)$ such that
$\operatorname{SR}_{\mathbb{F}_{q}}(T)\leq C\operatorname{SR}_{\overline{\mathbb{F}}_{q}}\left(T^{\overline{\mathbb{F}}_{q}}\right).$
Proof.
If $\mathbb{K}$ is a perfect infinite field. In the proof of Lemma 5.2, we may choose $c_{1}=d$ and $c_{2}=1/d$. Thus we have $C=d^{2}$ in Lemma 5.2. If $q$ is a power of $p$, then we have
$\operatorname{SR}_{\mathbb{K}}(T)\leq\operatorname{SR}_{\mathbb{F}_{p}}(T)\leq C(p,d)\operatorname{SR}_{\overline{\mathbb{F}}_{p}}\left(T^{\overline{\mathbb{F}}_{p}}\right)=C(p,d)\operatorname{SR}_{\overline{\mathbb{K}}}\left(T^{\overline{\mathbb{K}}}\right),$
where the first inequality follows from the definition, the second inequality is obtained by Lemma 5.2 and the equality is the consequence of the fact that $\overline{\mathbb{K}}=\overline{\mathbb{F}}_{p}$. ∎
Remark 5.4.
It is proved in [14] that
$\frac{2\operatorname{SR}_{\mathbb{K}}(T)}{d}\leq\operatorname{R}^{G}_{\mathbb{K}}(T)=\operatorname{R}^{G}_{\overline{\mathbb{K}}}(T)\leq\operatorname{SR}_{\overline{\mathbb{K}}}(T),$
where $\operatorname{R}^{G}_{\mathbb{K}}(T)$ is the G-stable rank of $T$. This leads to the stability of the slice rank under field extensions with a better constant:
$\operatorname{SR}_{\mathbb{K}}(T)\leq\frac{d}{2}\operatorname{SR}_{\overline{\mathbb{K}}}\left(T^{\overline{\mathbb{K}}}\right).$ (20)
We recall that for any positive integers $t\leq s,m_{1},\dots,m_{s}$, there is a natural injective map $\tau$ defined by
[…]
where $\mathbb{K}[\mathsf{V}]$ denotes the ring of polynomials on a vector space $\mathsf{V}$. The image of this injection consists of all multilinear polynomials on $(\mathbb{K}^{m_{1}})^{\ast}\oplus\cdots\oplus(\mathbb{K}^{m_{s}})^{\ast}$. The following inequality is stated in [28, Theorem 5] with a sketched proof involving several lemmas. We provide below a shorter proof.
Lemma 5.5.
Let $\mathbb{K}$ be a field and let $T\in\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$ be a tensor. Then $\operatorname{GR}_{\mathbb{K}}(T)\leq\operatorname{PR}_{\mathbb{K}}(T)$.
Proof.
Since $\operatorname{GR}_{\mathbb{K}}(T)=\operatorname{GR}_{\overline{\mathbb{K}}}\left(T^{\overline{\mathbb{K}}}\right)$ and $\operatorname{PR}_{\overline{\mathbb{K}}}\left(T^{\overline{\mathbb{K}}}\right)\leq\operatorname{PR}_{\mathbb{K}}(T)$, we may assume that $\mathbb{K}$ is algebraically closed. We define an $n_{d}$-dimensional subspace $\mathsf{W}\coloneqq\langle T,(\mathbb{K}^{n_{d}})^{\ast}\rangle$ in $\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d-1}}$. Denote $\mathsf{V}\coloneqq(\mathbb{K}^{n_{1}})^{\ast}\oplus\cdots\oplus(\mathbb{K}^{n_{d-1}})^{\ast}$. Let $\tau$ be the injection defined in (21) and let $J$ be the ideal generated by $\tau(\mathsf{W})$ in $\mathbb{K}[\mathsf{V}]$. Obviously, the variety defined by $J$ coincides with $Z_{d}(T)$ defined in (4) and $\operatorname{GR}_{\mathbb{K}}(T)=\operatorname{codim}J$.
Suppose that $T=\sum_{i=1}^{r}a_{i}\otimes b_{i}$ is a partition rank decomposition of $T$ where $r\coloneqq\operatorname{PR}_{\mathbb{K}}(T)$ and $a_{i}$ dose not involve the $d$-th factor for each $1\leq i\leq r$. Thus, we have $r$ polynomials:
$f_{i}\coloneqq\tau(a_{i})\in\mathbb{K}[\mathsf{V}],1\leq i\leq r.$
Since $\mathsf{W}=\langle T,(\mathbb{K}^{n_{d}})^{\ast}\rangle=\sum_{i=1}^{r}a_{i}\otimes\langle b_{i},(\mathbb{K}^{n_{d}})^{\ast}\rangle$ and $J$ is generated by $\mathsf{W}$, we may derive $J\subseteq(f_{1},\dots,f_{r})$. According to [15, Theorem 10.2], we conclude that $\operatorname{codim}J\leq r$. ∎
In the following, we establish the equivalence among four conjectures concerning the partition rank.
Theorem 5.6 (Equivalence of conjectures).
Let $d$ be a positive integer and let $q$ be a power of $p$. The following are equivalent:
  1. (a)
    Partition rank v.s. analytic rank conjecture [21, 33, 1]: there exists a constant $C_{1}\coloneqq C_{1}(d,q)>0$ such that
    $\operatorname{PR}_{\mathbb{F}_{q}}(T)\leq C_{1}\operatorname{AR}_{\mathbb{F}_{q}}(T)$
    for any $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$.
  2. (b)
    Partition rank v.s. geometric rank conjecture [37, 1, 9]: there exists a constant $C_{2}\coloneqq C_{2}(d,q)>0$ such that
    $\operatorname{PR}_{\mathbb{F}_{q}}(T)\leq C_{2}\operatorname{GR}_{\mathbb{F}_{q}}(T)$
    for any $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$.
  3. (c)
    Asymptotic direct sum conjecture for partition rank: there exists a constant $C_{3}\coloneqq C_{3}(d,q)>0$ such that
    $\operatorname{PR}_{\mathbb{F}_{q}}(T)\leq C_{3}\limsup_{n\to\infty}\frac{\operatorname{PR}_{\mathbb{F}_{q}}\left(T^{\oplus n}\right)}{n}$
    for any $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$.
  4. (d)
    Stability conjecture for partition rank [1]: there exists a constant $C_{4}\coloneqq C_{4}(d,q)>0$ such that
    $\operatorname{PR}_{\mathbb{F}_{q}}(T)\leq C_{4}\operatorname{PR}_{\overline{\mathbb{F}}_{q}}\left(T^{\overline{\mathbb{F}}_{q}}\right)$
    for any $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$.
Moreover, (a)–(d) hold for any powers of $p$ if and only if they hold for $p$.
Proof.
The implication (a) $\implies$ (b) follows from Proposition 4.4. We notice that by definition,
$\operatorname{GR}_{\mathbb{F}_{q}}(T\oplus S)=\operatorname{GR}_{\mathbb{F}_{q}}(T)+\operatorname{GR}_{\mathbb{F}_{q}}(S).$
Thus, we have $\operatorname{PR}_{\mathbb{F}_{q}}\left(T^{\oplus n}\right)\geq\operatorname{GR}_{\mathbb{F}_{q}}\left(T^{\oplus n}\right)=n\operatorname{GR}_{\mathbb{F}_{q}}(T)$ by Lemma 5.5. This proves (b) $\implies$ (c). If (c) holds, then (d) also holds by Lemma 5.2. The implication (d) $\implies$ (a) is obtained by combining [9, Corollary 3] and the proof of [1, Theorem 1.13].
One direction of the ’moreover’ part is trivial. To prove the other direction, it is sufficient to prove that (c) holds for $q=p^{l}$ for any $l\in\mathbb{N}$ as long as it holds for $p$. Given a tensor $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$, Theorem 3.2 and Lemma 5.1 indicate the existence of a constant $C(d,p)$ such that $\operatorname{PR}_{\mathbb{F}_{p}}(T^{\oplus n})\leq C(d,p)l\operatorname{PR}_{\mathbb{F}_{q}}(T^{\oplus n})$. Since (c) holds for $p$, we also have $c(d,p)n\operatorname{PR}_{\mathbb{F}_{p}}(T)\leq\operatorname{PR}_{\mathbb{F}_{p}}(T^{\oplus n})$. We obtain
$\dfrac{c(d,p)\operatorname{PR}_{\mathbb{F}_{q}}(T)}{C(d,p)l}\leq\dfrac{c(d,p)\operatorname{PR}_{\mathbb{F}_{p}}(T)}{C(d,p)l}\leq\dfrac{\operatorname{PR}_{\mathbb{F}_{p}}(T^{\oplus n})}{C(d,p)ln}\leq\dfrac{\operatorname{PR}_{\mathbb{F}_{q}}(T^{\oplus n})}{n},$
where the first inequality follows from Lemma 5.1. Hence (c) for $q$ is verified by taking $C^{\prime}(d,q)\coloneqq C(d,p)l/c(d,p)$. ∎
We remark that the dependence of constants $C_{1},\dots,C_{4}$ on $q$ in Theorem 5.6 can be removed by a similar argument [35]. However, Theorem 5.6 and its base field independence version are actually parallel, in the sense that one can not be obtained from the other.
Next we consider the slice rank of a linear subspace of tensors. Given a linear subspace $\mathsf{W}\subseteq\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$ over a field $\mathbb{K}$, the slice rank of $\mathsf{W}$ is defined as
[…]
Here $\langle\mathsf{W},\mathsf{U}_{1}\otimes\cdots\otimes\mathsf{U}_{d}\rangle=0$ means that $\langle T,L\rangle=0$ for each pair $(T,L)\in\mathsf{W}\times\left(\mathsf{U}_{1}\otimes\cdots\otimes\mathsf{U}_{d}\right)$. In the sequel, we denote
$\mathsf{V}_{i}\coloneqq\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{i-1}}\otimes\mathbb{K}^{n_{i+1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}},$ (23)
for each $1\leq i\leq d$.
Lemma 5.7 (Characterization of $\operatorname{SR}_{\mathbb{K}}(\mathsf{W})$).
Let $\mathbb{K}$ be a field. For a linear subspace $\mathsf{W}\subseteq\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$, $\operatorname{SR}_{\mathbb{K}}(\mathsf{W})$ is the minimal positive integer $r$ to ensure the existence of $r_{i}\in\mathbb{N}$ and $a_{i1},\dots,a_{ir_{i}}\in\mathbb{K}^{n_{i}}$ for $1\leq i\leq d$, such that $r=\sum_{i=1}^{d}r_{i}$ and every $T\in\mathsf{W}$ admits a decomposition
$T=\sum_{i=1}^{d}\sum_{j=1}^{r_{i}}a_{ij}\otimes T_{ij}$ (24)
for some $T_{i1},\dots,T_{ir_{i}}\in\mathsf{V}_{i}$, $1\leq i\leq d$.
Proof.
We denote by $s$ the minimal integer such that (24) holds. We prove that $s=t\coloneqq\operatorname{SR}_{\mathbb{K}}(\mathsf{W})$. If $\langle\mathsf{W},\mathsf{U}_{1}\otimes\cdots\otimes\mathsf{U}_{d}\rangle=0$, then $\mathsf{W}\subseteq\sum_{i=1}^{d}\mathsf{U}_{i}^{\perp}\otimes\mathsf{V}_{i}$ where $\mathsf{U}_{i}^{\perp}\coloneqq((\mathbb{K}^{n_{i}})^{\ast}/\mathsf{U}_{i})^{\ast}$. Hence any $T\in\mathsf{W}$ admits a decomposition (24) if we choose $r_{i}=\dim\mathsf{U}_{i}^{\perp}$ and let $\{a_{i1},\dots,a_{ir_{i}}\}$ be a basis of $\mathsf{U}_{i}^{\perp}$ for $1\leq i\leq d$. Thus we have $t=\sum_{i=1}^{d}\dim\mathsf{U}_{i}^{\perp}=\sum_{i=1}^{d}r_{i}\geq s$.
Conversely, suppose that $r_{i}$’s and $a_{ij}$’s are given so that $s=\sum_{i=1}^{d}r_{i}$ and (24) holds for each $T\in\mathsf{W}$. We define
$\mathsf{U}_{i}\coloneqq\left(\operatorname{span}\{a_{i1},\dots,a_{ir_{i}}\}\right)^{\perp}\coloneqq\left(\mathbb{K}^{n_{i}}/\operatorname{span}\{a_{i1},\dots,a_{ir_{i}}\}\right)^{\ast}\subseteq(\mathbb{K}^{n_{i}})^{\ast},\quad 1\leq i\leq d.$
Then clearly we have $\langle\mathsf{W},\mathsf{U}_{1}\otimes\cdots\otimes\mathsf{U}_{d}\rangle=0$ and this implies $s=\sum_{i=1}^{d}r_{i}\geq t$. ∎
Lemma 5.8.
For any field $\mathbb{K}$ and linear subspace $\mathsf{W}\subseteq\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$, there exist $n_{d+1},n_{d+2}\in\mathbb{N}$ and $T_{\mathsf{W}}\in\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d+2}}$ such that $\operatorname{SR}_{\mathbb{K}}(\mathsf{W})=\operatorname{SR}_{\mathbb{K}}(T_{\mathsf{W}})$. .
Proof.
[…]
$T_{\mathsf{W}}\coloneqq\sum_{j=1}^{n}T_{j}\otimes\left(\sum_{i=1}^{m}e_{ij}\otimes e_{ij}\right)\in\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}\otimes\mathbb{K}^{mn}\otimes\mathbb{K}^{mn}.$
By Lemma 5.7, it is clear that $\operatorname{SR}_{\mathbb{K}}(T_{\mathsf{W}})\leq\operatorname{SR}_{\mathbb{K}}(\mathsf{W})$. Thus, it suffices to prove $\operatorname{SR}_{\mathbb{K}}(T_{\mathsf{W}})\geq\operatorname{SR}_{\mathbb{K}}(\mathsf{W})$.
Let $T_{\mathsf{W}}=\sum_{i=1}^{d+2}\sum_{j=1}^{r_{i}}a_{ij}\otimes S_{ij}$ be a slice rank decomposition of $T_{\mathsf{W}}$, where $a_{ij}\in\mathbb{K}^{n_{i}}$ and $S_{ij}\in\mathsf{V}_{i}$ for $21\leq i\leq d+$ 2 and $1\leq j\leq r_{i}$. Here $\mathsf{V}_{i}$ is defined as in (23). We denote
$\mathsf{W}^{\prime}\coloneqq\sum_{i=1}^{d}\sum_{j=1}^{r_{i}}\left\langle a_{ij}\otimes S_{ij},(\mathbb{K}^{mn})^{\ast}\otimes(\mathbb{K}^{mn})^{\ast}\right\rangle\subseteq\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}.$
We observe that each $L\in\mathsf{W}^{\prime}$ is a linear combination of $a_{ij}\otimes S^{\prime}_{ij}$ for some
[…]
Without loss of generality, we may assume $T_{1},\cdots,T_{l}\in W^{\prime}$ and $T_{l+1},\cdots,T_{n}\notin W^{\prime}$ for some $0\leq l\leq n$. Therefore we have
$\operatorname{SR}_{\mathbb{K}}(\mathsf{W})\leq\sum_{i=1}^{d}r_{i}+\sum_{s=l+1}^{n}\max\{n_{i}\}_{i=1}^{d}\leq\sum_{i=1}^{d}r_{i}+(n-l)m.$
In the following, we regard $a_{ij}$’s (resp. $T_{k}$’s) as linear (resp. multilinear) polynomials. Since $T_{l+1},\dots,T_{n}\not\in\mathsf{W}^{\prime}$, we conclude that $T_{l+1}\cdots T_{n}\notin J$, where $J$ is the prime ideal generated by linear forms $a_{ij},1\leq i\leq d,1\leq j\leq r_{i}$. By Hilbert Nullstellensatz, there exist $v_{i}\in\overline{\mathbb{K}}^{n_{i}}$, $1\leq i\leq d$, such that
[…]
Therefore, we obtain
$T_{W}(v_{1},\cdots,v_{d})$ $=\sum_{i=1}^{d+2}\sum_{j=1}^{r_{i}}(a_{ij}\otimes S_{ij})(v_{1},\dots,v_{d})$
$=\sum_{j=1}^{r_{d+1}}a_{d+1,j}\otimes S_{d+1,j}(v_{1},\dots,v_{d})+\sum_{j=1}^{r_{d+2}}a_{d+2,j}\otimes S_{d+2,j}(v_{1},\dots,v_{d})$ (25)
On the other hand, by the construction of $T_{\mathsf{W}}$, we also have
$T_{W}(v_{1},\cdots,v_{d})=\sum_{s=1}^{n}T_{s}(v_{1},\dots,v_{d})\left(\sum_{i=1}^{m}e_{is}\otimes e_{is}\right)=\sum_{s=l+1}^{n}T_{s}(v_{1},\dots,v_{d})\left(\sum_{i=1}^{m}e_{is}\otimes e_{is}\right).$ (26)
Both (25) and (26) are decompositions of the matrix $T_{W}(v_{1},\cdots,v_{d})\in\overline{\mathbb{K}}^{mn\times mn}$ into the sum of rank one matrices. Moreover, (26) is clearly a rank decomposition. This implies that $(n-l)m\leq r_{d+1}+r_{d+2}$ and
$\operatorname{SR}_{\mathbb{K}}(\mathsf{W})\leq\sum_{i=1}^{d}r_{i}+(n-l)m\leq\sum_{i=1}^{d+2}r_{i}=\operatorname{SR}_{\mathbb{K}}(T_{\mathsf{W}}).\qed$
A direct application of Lemmas 5.8 and 2.1 implies the additivity of the slice rank of linear subspaces with respect to the direct sum.
Proposition 5.9 (Additivity of slice rank of subspaces).
For any field $\mathbb{K}$ and linear subspaces $\mathsf{W}_{1}\subseteq\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$, $\mathsf{W}_{2}\subseteq\mathbb{K}^{m_{1}}\otimes\cdots\otimes\mathbb{K}^{m_{d}}$, we have
$\operatorname{SR}_{\mathbb{K}}(\mathsf{W}_{1}\oplus\mathsf{W}_{2})=\operatorname{SR}_{\mathbb{K}}(\mathsf{W}_{1})+\operatorname{SR}_{\mathbb{K}}(\mathsf{W}_{2}).$
Conjecture 1.5 can be easily confirmed by Lemma 5.8.
Theorem 5.10 (Stability of slice rank of subspaces I).
Let $\mathbb{K}$ be a field. For any positive integer $d$ and linear subspace $\mathsf{W}\subseteq\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$, we have
$\operatorname{SR}_{\mathbb{K}}(\mathsf{W})\leq\left(\frac{d}{2}+1\right)\operatorname{SR}_{\overline{\mathbb{K}}}\left(\mathsf{W}\otimes\overline{\mathbb{K}}\right).$
Proof.
According to the proof of Lemma 5.8, there is a tensor $T_{\mathsf{W}}\in\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d+2}}$ such that $\operatorname{SR}_{\mathbb{K}}(\mathsf{W})=\operatorname{SR}_{\mathbb{K}}(T_{\mathsf{W}})$ and $\operatorname{SR}_{\overline{\mathbb{K}}}(\mathsf{W}\otimes\overline{\mathbb{K}})=\operatorname{SR}_{\overline{\mathbb{K}}}\left(T_{\mathsf{W}}^{\overline{\mathbb{K}}}\right)$. Now the desired inequality follows immediately from (20). ∎
We remark that Theorem 5.10 is proved for $d=2$ in [1] by a different approach. To conclude this section, we define for each $\mathsf{W}\subseteq\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$ and positive integer $1\leq k\leq\dim\mathsf{W}$ the following quantity:
$\operatorname{SR}_{k,\mathbb{K}}(\mathsf{W})\coloneqq\max\{\operatorname{SR}_{\mathbb{K}}(\mathsf{V}):\mathsf{V}\subseteq\mathsf{W},\;\dim\mathsf{V}=k\}.$ (27)
We will prove below that for $d=2$, $\operatorname{SR}_{k,\mathbb{K}}(\mathsf{W})$ lies between constant multiples of $\operatorname{SR}_{\mathbb{K}}(\mathsf{W})$ and $\operatorname{SR}_{\overline{\mathbb{K}}}(\mathsf{W}\otimes\overline{\mathbb{K}})$.
Lemma 5.11.
For any field $\mathbb{K}$ and linear subspace $\mathsf{W}\subseteq\mathbb{K}^{n_{1}\times n_{2}}$ and integer $1\leq k\leq\dim\mathsf{W}$, we have
$\operatorname{SR}_{\mathbb{K}}(W)\leq 2\operatorname{SR}_{k,\mathbb{K}}(\mathsf{W}).$
Proof.
The argument is similar to that in the proof of [1, Theorem 1.17], thus we omit it. ∎
Proposition 5.12 (Stability of slice rank of subspaces II).
Let $\mathbb{K}$ be a field. For any linear subspace $\mathsf{W}\subseteq\mathbb{K}^{n_{1}\times n_{2}}$ and integer $1\leq k\leq\dim\mathsf{W}$, we have
$\frac{1}{2}\operatorname{SR}_{\mathbb{K}}(\mathsf{W})\leq\operatorname{SR}_{k,\mathbb{K}}(\mathsf{W})\leq 2\operatorname{SR}_{k,\overline{\mathbb{K}}}(\mathsf{W}\otimes\overline{\mathbb{K}})\leq 2\operatorname{SR}_{\overline{\mathbb{K}}}(\mathsf{W}\otimes\overline{\mathbb{K}}).$
Proof.
The first inequality is obtained by Lemma 5.11 and the last is clear by definition. For the second inequality, let $\mathsf{V}$ be a $k$-dimensional subspace of $\mathsf{W}$ such that $\operatorname{SR}_{k,\mathbb{K}}(\mathsf{W})=\operatorname{SR}_{\mathbb{K}}(\mathsf{V})$. Then Theorem 5.10 implies
$\operatorname{SR}_{k,\mathbb{K}}(\mathsf{W})=\operatorname{SR}_{\mathbb{K}}(\mathsf{V})\leq 2\operatorname{SR}_{\overline{\mathbb{K}}}(\mathsf{V}\otimes\overline{\mathbb{K}})\leq 2\operatorname{SR}_{k,\overline{\mathbb{K}}}(\mathsf{W}\otimes\overline{\mathbb{K}}).\qed$
It is natural to expect that for any integer $d\geq 3$, there is also a constant $C\coloneqq C(d)>0$ such that
$\operatorname{SR}_{\mathbb{K}}(\mathsf{W})\leq C\operatorname{SR}_{k,\mathbb{K}}(\mathsf{W}).$ (28)
If (28) holds, then our proof of Proposition 5.12 can be generalized for $d\geq 3$. Unfortunately, this is not always the case by the proposition that follows.
Proposition 5.13.
For $d\geq 3$ and $k=1$, there is no constant $C\coloneqq C(d)>0$ such that (28) holds.
Proof.
We only prove the non-existence of $C$ for $d=3$, as the proof for larger $d$ is the same. Let $n$ be a positive integer. We consider tensors
$T_{s}\coloneqq I_{n}\otimes e_{k}\in\mathbb{K}^{n}\otimes\mathbb{K}^{n}\otimes\mathbb{K}^{n},\quad 1\leq s\leq n,$
where $I_{n}$ is the $n\times n$ identity matrix and $\{e_{1},\dots,e_{n}\}$ is a basis of $\mathbb{K}^{n}$. Denote $\mathsf{W}\coloneqq\operatorname{span}\{T_{1},\dots,T_{n}\}$. It is clear that $\operatorname{SR}_{1,\mathbb{K}}(\mathsf{W})=1$.
Let $1\leq r_{1},r_{2},r_{3}\leq n$ be positive integers such that $r_{1}+r_{2}+r_{3}=\operatorname{SR}_{\mathbb{K}}(\mathsf{W})$ and
$T_{s}\in\mathsf{U}_{1}\otimes\mathbb{K}^{n}\otimes\mathbb{K}^{n}+\mathbb{K}^{n}\otimes\mathsf{U}_{2}\otimes\mathbb{K}^{n}+\mathbb{K}^{n}\otimes\mathbb{K}^{n}\otimes\mathsf{U}_{3},\quad 1\leq s\leq n,$
for some $\mathsf{U}_{i}\subseteq\mathbb{K}^{n}$ of dimension $r_{i}$, $i=1,2,3$. If $r_{3}<n$, then there exists $f\in\mathsf{U}_{3}^{\perp}\coloneqq(\mathbb{K}^{n}/\mathsf{U}_{1})^{\ast}$ such that $f(e_{k_{0}})=1$. This implies
$I_{n}=\langle T_{k_{0}},f\rangle\in\mathsf{U}_{1}\otimes\mathbb{K}^{n}+\mathbb{K}^{n}\otimes\mathsf{U}_{2}.$
Hence we obtain $\operatorname{SR}_{\mathbb{K}}(\mathsf{W})\geq r_{1}+r_{2}\geq\operatorname{R}_{\mathbb{K}}(I_{n})=n$. If $r_{3}=n$, then we also have $\operatorname{SR}_{\mathbb{K}}(\mathsf{W})\geq n$. Since $n$ can be arbitrarily large and $\operatorname{SR}_{1,\mathbb{K}}(\mathsf{W})=1$, there is no constant $C>0$ such that $\operatorname{SR}_{\mathbb{K}}(\mathsf{W})\leq C\operatorname{SR}_{1,\mathbb{K}}(\mathsf{W})$. ∎
We notice that (28) holds for any $d$ when $k=\dim\mathsf{W}$. Thus, it is interesting to study the critical value $k_{0}$ such that (28) holds for $k\geq k_{0}$. According to Theorem 5.10 and Proposition 5.12, it is also reasonable to expect that $\operatorname{SR}_{k,\mathbb{K}}(\mathsf{W})$ is stable under field extensions as well.
Conjecture 5.14 (Stability of slice rank of subspaces III).
There is a constant $C\coloneqq C(d)>0$ such that for any linear subspace $\mathsf{W}\subseteq\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$ and positive integer $k\leq\mathsf{W}$, we have
$\operatorname{SR}_{k,\overline{\mathbb{K}}}(W\otimes\overline{\mathbb{K}})\leq C\operatorname{SR}_{k,\mathbb{K}}(\mathsf{W}).$
Acknowledgement
We would like to thank Daniel Zhu and Guy Moshkovitz for their helpful comments on the previous version of this manuscript. With their suggestions, we fill gaps in the proofs of Lemma 3.5 and Theorem 3.7.
Appendix A Algebraic function field
We provide a brief introduction to the theory of algebraic function fields. Interested readers are referred to [40] for more details. An algebraic function field of one variable over $\mathbb{K}$ is a field extension $\mathbb{F}/\mathbb{K}$ such that there is a transcendental element $f\in\mathbb{F}$ and $[\mathbb{F}:\mathbb{K}(f)]<\infty$. We recall that any valuation ring $(\mathcal{O},\mathfrak{m},v_{\mathfrak{m}})$ such that $\mathbb{K}\subsetneq\mathcal{O}\subsetneq\mathbb{F}$ and $F(\mathcal{O})=\mathbb{F}$ is a discrete valuation ring (DVR). Here $\mathfrak{m}$ is the maximal ideal of $\mathcal{O}$, $F(\mathcal{O})$ is the field of fractions of $\mathcal{O}$ and $v_{\mathfrak{m}}$ is the valuation of $F(\mathcal{O})$. The maximal ideal $\mathfrak{m}$ of $\mathcal{O}$ is called a place of $\mathbb{F}/\mathbb{K}$. The degree of $\mathfrak{m}$ is defined as $\deg(\mathfrak{m})\coloneqq\dim_{\mathbb{K}}(\mathcal{O}/\mathfrak{m})$.
Let $\operatorname{Div}(\mathbb{F}/\mathbb{K})$ be the free abelian group generated by places of $\mathbb{F}/\mathbb{K}$. An element in $\operatorname{Div}(\mathbb{F}/\mathbb{K})$ is called a divisor and $\operatorname{Div}(\mathbb{F}/\mathbb{K})$ is called the divisor group. For each $D\in\sum_{\mathfrak{m}}n_{\mathfrak{m}}\mathfrak{m}$ where $n_{\mathfrak{m}}=0$ for all but finitely many $\mathfrak{m}$’s, we define the degree of $D$ by $\deg(D)\coloneqq\sum_{\mathfrak{m}}n_{\mathfrak{m}}\deg(\mathfrak{m})$. If $n_{\mathfrak{m}}\geq 0$ for each place $\mathfrak{m}$, we denote $D\geq 0$. Given an element $f\in\mathbb{F}$, the principal divisor generated by $f$ is $(f)\coloneqq\sum_{\mathfrak{m}}v_{\mathfrak{m}}(f)\mathfrak{m}$.
For each divisor $D$, the Riemann-Roch space is defined as
$\mathscr{L}(D)\coloneqq\left\{f\in\mathbb{F}:(f)+D\geq 0\right\}\cup\left\{0\right\}$
We notice that $\mathscr{L}(D)$ is a finite dimensional $\mathbb{K}$-subspace of $\mathbb{F}$ and denote $\ell(D)\coloneqq\dim_{\mathbb{K}}\mathscr{L}(D)$. The genus of $\mathbb{F}/\mathbb{K}$ is $g\coloneqq\max_{D\in\operatorname{Div}(\mathbb{F}/\mathbb{K})}\left\{\deg(D)-\ell(D)+1\right\}$. For ease of reference, we record below some basic lemmas that will be needed in the paper.
Lemma A.1.
[40, Theorem 1.5.17] Let $g$ be the genus of $\mathbb{F}/\mathbb{K}$. For each $D\in\operatorname{Div}(\mathbb{F}/\mathbb{K})$ with $\deg(D)\geq 2g-1$, we have $\ell(D)=\deg(D)+1-g$.
Lemma A.2.
[40, Theorem 1.4.11] All principal divisors have degree 00.
Lemma A.3 (Weak approximation).
[40, Theorem 1.3.1]] Let $\mathbb{F}/\mathbb{K}$ be a function field. Given distinct places $\mathfrak{m}_{1},\dots,\mathfrak{m}_{n}$ of $\mathbb{F}/\mathbb{K}$, $f_{1},\cdots,f_{n}\in\mathbb{F}$ and $r_{1},\dots,r_{n}\in\mathbb{Z}$, there exists some $f\in\mathbb{F}$ such that $v_{\mathfrak{m}_{i}}(f-f_{i})=r_{i}$ for each $1\leq i\leq n$.
Lemma A.4.
[40, Corollary 1.4.12] For any $D\in\operatorname{Div}(\mathbb{F}/\mathbb{K})$ with $\deg(D)<0$, we have $\ell(D)=0$.
A divisor $D\in\operatorname{Div}(\mathbb{F}/\mathbb{K})$ is non-special if $\ell(D)=\deg(D)+1-g$, otherwise it is special.
Lemma A.5.
[2, Lemma 2.1] Let $q$ be a prime power. Let $\mathbb{F}/\mathbb{F}_{q}$ be an algebraic function field of genus $g$ containing at least $g+1$ places of degree one. Denote by $P_{1}(F/\mathbb{F}_{q})$ the set of places of degree one. For any $S\subseteq P_{1}(\mathbb{F}/\mathbb{F}_{q})$ with $|S|\geq g+1$, there exists a non-special divisor $D$ such that $\deg(D)=g-1$ and $supp(D)\subseteq S$. Here $supp(D)$ is the set of place with non-zero coefficient in $D$.
Lemma A.6.
[17, Theorem] For each positive integer $k$, there exists an algebraic function field $\mathbb{F}/\mathbb{F}_{q^{2}}$ whose genus is
$g_{k}\coloneqq\begin{cases}q^{k}+q^{k-1}-q^{(k+1)/2}-2q^{(k-1)/2}+1&\text{if $k$ is odd}\\ q^{k}+q^{k-1}-\frac{1}{2}q^{k/2+1}-\frac{3}{2}q^{k/2}-q^{k/2-1}+1&\text{otherwise}\end{cases}.$ (29)
Moreover, if $k\geq 3$, then $\mathbb{F}$ admits at least
$N_{k}\coloneqq(q^{2}-1)q^{k-1}+2q\geq(q-1)g_{k}$ (30)
degree one places.
Lemma A.7.
[40, Corollary 5.2.10] Any algebraic function field $\mathbb{F}/\mathbb{F}_{q}$ of genus $g$ admits at least one place of degree $r$ for each $r\geq 4g+3$.
Appendix B Number of rational points
Theorem B.1.
[19, Theorem 11.1] Let $X$ be an $n$-dimensional irreducible projective subvariety of $\mathbb{P}_{\overline{\mathbb{F}}_{q}}^{N}$ defined over $\mathbb{F}_{q}$. Denote $d\coloneqq\deg(X)$. There is a constant number $C\coloneqq C(X)>0$ such that
$\left\lvert|X(\mathbb{F}_{q})|-\frac{q^{n+1}-1}{q-1}\right\rvert\leq(d-1)(d-2)q^{n-\frac{1}{2}}+Cq^{n-1}.$